# Reference map technique for Lagrangian exploration of coherent structures

### v1

@@ Abstract
We explore the application of the reference map technique, originally developed for Eulerian simulation of solid mechanics, in Lagrangian kinematics of fluid flows. Unlike traditional methods based on explicit particle tracking, the reference map facilitates calculation of flow maps and gradients without the need for particles. This is achieved through an Eulerian update of the reference map, which records the take-off positions of fluid particles. Using the reference map, measures of Lagrangian fluid-element deformation can be computed solely from Eulerian data, such as displacement, deformation gradient, Cauchy-Green tensor, and Finite-Time Lyapunov Exponent (FTLE) fields. We first demonstrate the accuracy of FTLE calculations based on the reference map against the standard particle-based approach in a 2D Taylor-Green vortex. Then, we apply it to turbulent channel flow at $Re_{\tau}=180$ where Lagrangian coherent structures identified as ridges of the backward-time FTLE are found to bound vortical regions of flow, consistent with Eulerian coherent structures from Q-criterion. The reference map also proves suitable for material surface tracking despite not explicitly tracking particles. This capability can provide valuable insights into the Lagrangian landscape of turbulent momentum transport, complementing Eulerian velocity-field analysis. The evolution of initially wall-normal material surfaces in the viscous sublayer, buffer layer, and log layer sheds light on the Reynolds stress-generating events from a Lagrangian perspective. Eliminating the need for tracking numerous particles, the reference map approach offers a convenient and promising avenue for future investigations into Lagrangian kinematics and dynamics of fluid flows.
@@ 1 Introduction
Fluid flows are often characterized by dynamic behaviors of intricate spatial patterns. Notable examples include bacteria-induced active-matter chaos (Ran et al., 2021; Urzay et al., 2017), a forest of vigorous turbulent eddies in wing boundary layers (Goc et al., 2021), and the great red spot of Jupiter (Marcus, 1993). While visually captivating, the complex interaction among these spatial patterns can sometimes obscure straightforward comprehension of the dynamically important mechanisms governing them. The concept of coherent structures has proven integral to identifying these organized features and furthering the analysis and modeling of their dynamical impact on the flow. Coherent structures, broadly defined, represent persistent patterns or organized motions that stand out against the background flow. In a more specific sense, Hussain (1986) defined coherent structures in turbulent flows as ”a connected turbulent fluid mass with instantaneously phase-correlated vorticity over its spatial extent.” Meanwhile, Haller (2015) offered a definition strongly rooted in Lagrangian/material kinematics and dynamical systems theory, describing them as ”special surfaces of fluid trajectories that organize the rest of the flow into ordered patterns.” Despite differences in the perspectives and approaches taken to define them, it is perhaps unanimously agreed that coherent structures serve as the engines driving the complex interplay of momentum, heat, and scalar transport at various regimes.
@@
Recognizing the crucial role coherent structures play in fluid dynamics, the fluids community has dedicated considerable attention to developing methods for elucidating and understanding these organized features. A widely used class of Eulerian methods, particularly within the turbulence community, involves vortex identification schemes based on the invariants of the velocity gradient tensor or its symmetric/antisymmetric decomposition. The well-known $Q$ criterion (Hunt et al., 1988) extracts vortical structures from the second invariant of the velocity gradient tensor $\nabla u$, defined as $Q=-\frac{1}{2}\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{j}}{\partial x_{i}}=\frac{1}{2}(||\Omega||^{2}-||S||^{2})$, where $\Omega$ and $S$ are the rate-of-rotation and rate-of-strain tensors, respectively. $Q$ is regarded as the strength of the rotational component of the velocity field relative to the shear, and isosurfaces of $Q$ at a selected positive threshold value are used as candidate coherent structures. In response to potential inaccuracies of $Q$ criterion when vortices are subjected to strong external strain, Jeong & Hussain (1995) proposed the $\lambda_{2}$ criterion. The $\lambda_{2}$ criterion imposes the condition that $\lambda_{2}<0$, where $\lambda_{2}$ is the median of three eigenvalues of $S^{2}+\Omega^{2}$. This requirement is based on the principles of Galilean invariance and the presence of net vorticity, which is correlated with a local pressure minimum at the vortex core. Zhou et al. (1999) proposed a detection scheme based on complex eigenvalues of $\nabla u$. From the conjugate pair of the complex eigenvalues $\lambda_{cr}\pm i\lambda_{ci}$ of $\nabla u$, the imaginary part $\lambda_{ci}$ was shown to represent the local swirling strength of the vortex. In an extension of this work, Chakraborty et al. (2005) suggested an additional constraint $\lambda_{cr}/\lambda_{ci}<\delta$ to ensure the compactness of the spiral material orbits. The aforementioned detection methods are computed from the instantaneous Eulerian velocity/tensor fields, making their application to spatial velocity data obtained from numerical simulations or experiments straightforward. They have proven effective in excluding regions of strong shear without net swirling motion. These methods necessitate users to prescribe arbitrary thresholds for isosurface detection. As demonstrated by Pierce et al. (2013), all the aforementioned schemes produce nearly identical landscapes of turbulent eddies with an appropriate selection of threshold values for each scheme.
@@
In contrast to Eulerian methods described in the previous paragraph, Lagrangian coherent structures (LCS) offer an alternative approach to coherent structure detection with an emphasis on coherence at the material (Lagrangian) level and objectivity. Haller (2005) demonstrated that the previously mentioned Eulerian diagnostics, while Galilean invariant, fail the test of objectivity, where the coherent structures identified should be invariant under Euclidean coordinate changes, including time-dependent frame rotation and translation. It’s essential to note that this does not diminish the utility of Eulerian diagnostics altogether. These approaches effectively reveal coherent patterns in the instantaneous velocity field, aiding in understanding vortex interaction/formation and momentum transport in wall-bounded flows (Jeong & Hussain, 1995; Zhou et al., 1999; Duguet et al., 2012; Motoori & Goto, 2019; Yao & Hussain, 2020; Chong et al., 1998). However, limitations may arise in the presence of frame rotation, and structural models built from the detected patterns may have constraints concerning objectivity. Haller recognized that objectivity can be ensured by basing coherent structure detection on fluid particle trajectories, which are fundamentally invariant Lagrangian features of a given flow. The representation of trajectories may vary in different reference frames, but the trajectories themselves remain invariant. The spatial patterns created by passively advected tracers then naturally embody the candidates for LCS. Haller (2015) explains LCS as ”the most repelling, attracting, and shearing material surfaces that form the skeleton of Lagrangian particle dynamics.” As such, LCS is expected to have direct connections to the motion of Lagrangian fluid elements.
@@ 2 Theoretical background
First, the flow map (also referred to as deformation mapping or motion map) $\bm{\chi}^{t}_{t_{0}}(\textbf{X},t)$ maps a fluid particle (tracer) located at X in the reference (or initial) configuration at the initial time $t_{0}$ to a point x in the current configuration of the continuum body at time $t$ (see figure 1).
$\textbf{x}=\bm{\chi}^{t}_{t_{0}}(\textbf{X},t).$ (1)
The deformation gradient (or Jacobian of the flow map) $\mathbf{F}^{t}_{t_{0}}(\textbf{X},t)$ is defined as
@@
which maps a material point x in the current configuration at time $t$ to its initial position X in the reference configuration. Hereinafter, this inverse flow map $\bm{\xi}$ is referred to as the reference map following Kamrin et al. (2012). The deformation gradient $\textbf{F}^{t}_{t_{0}}(\textbf{X},t)$, which is the Jacobian of the forward motion map at the initial position X, can be determined from the reference map.
$\left.\textbf{F}^{t}_{t_{0}}(\textbf{X},t)\right|_{\textbf{X}=\bm{\xi}(\textbf{x},t)}=\left(\frac{\partial\bm{\xi}(\textbf{x},t)}{\partial\textbf{x}}\right)^{-1}.$ (4)
Analogously, the inverse deformation gradient (or the Jacobian of the backward motion map at the current position x) is defined as
@@
The reference map framework was developed originally in the context of isothermal compressible and incompressible fluid-structure interaction (Kamrin et al., 2012; Valkov et al., 2015; Jain et al., 2019; Rycroft et al., 2020; Wang et al., 2022). Additionally, it has been employed in computational solid mechanics with non-trivial hyperelastic constitutive models undergoing inhomogeneous deformation (Kamrin et al., 2012). In this line of work, the Lagrangian solid equation recast in the Eulerian form, which resembles the Navier-Stokes equations, was solved using finite difference methods. The deformation gradient, determining the solid stress, was updated from the reference map. The key starting point was to identify x, the current position of material points, as the fixed grid points in Eulerian simulations or measurements. $\bm{\xi}(\textbf{x},t)$ then returns the initial position X of a Lagrangian material point that is arriving at the grid point x at time $t$. Since a material point labeled by X preserves its identity during the motion, the Lagrangian time derivative of the reference map is 0.
@@
In the present study, we employ the reference map to compute the Lagrangian kinematics of a fluid, where the reference map is updated through Eq.(6) in an Eulerian manner. That is, the reference map is passively advected on the same fixed Eulerian grid utilized in fluid dynamics simulations or from experimental velocity measurements. The deformation gradient is then determined entirely from the reference map, which can be further processed to compute Lagrangian coherent structures (LCSs). Additionally, the reference map provides a natural means for material surface tracking. Details of these two tasks are provided in the following subsections. Computational details associated with the discretization of Eq.(6) and initial/boundary conditions for the reference map are provided in Appendix A.
@@ 2.1 Computation of Lagrangian coherent structure (LCS)
    $\Lambda^{t_{0}}_{t}(\textbf{x})=\frac{1}{|t-t_{0}|}\text{ln}\sqrt{\lambda_{\text{max}}(\textbf{C}^{t_{0}}_{t}(\textbf{x}))}$ (12)
@@
Note that the backward-time FTLE is defined in the current configuration (x), which can be directly computed at the current position (grid points) using the reference map technique. Conversely, the forward-time FTLE is, in principle, defined in the space of the reference configuration X, which can be directly computed using a Lagrangian particle approach. However, the reference map technique can be used to compute the image, under the forward flow map, of the FTLE field in the current configuration. This can be shown by inverting Proposition 1 from Haller & Sapsis (2011) to find
@@
which shows that the largest forward-time FTLE can be obtained from the smallest backward-time FTLE defined at the current position x. Proof of this equation is provided in Appendix B.
@@ 2.2 Lagrangian tracking of material surface
In the original work by Kamrin et al. (2012) which introduced the reference map technique for fluid/structure interaction, the reference map was used primarily for computing the solid stress tensor, and the material interface was tracked by separately by integrating the level-set equation. Later, Jain et al. (2019) eliminated the need for the latter, observing that the reference map provides a direct mapping to the initial positions of the advected particles. This feature allows for convenient updates of the level-set function or material interface directly from the reference map. Adopting this concept, particularly from Eq. (25) in Jain et al. (2019), the reference map can be demonstrated to offer basic capabilities for Lagrangian tracking of material surfaces in purely fluid flows. For instance, in two-dimensional flows, consider a scalar field constructed from the reference map.

### v2

@@ Abstract
We explore the application of the reference map technique, originally developed for Eulerian simulation of solid mechanics, in Lagrangian kinematics of fluid flows. Unlike traditional methods based on explicit particle tracking, the reference map facilitates calculation of flow maps and gradients without the need for particles. This is achieved through an Eulerian update of the reference map, which records the take-off positions of fluid particles. Using the reference map, measures of Lagrangian fluid-element deformation can be computed solely from Eulerian data, such as displacement, deformation gradient, Cauchy-Green tensor, and Finite-Time Lyapunov Exponent (FTLE) fields. We first demonstrate the accuracy of FTLE calculations based on the reference map against the standard particle-based approach in a 2D Taylor-Green vortex. Then, we apply it to turbulent channel flow at $Re_{\tau}=180$ where Lagrangian coherent structures identified as ridges of the backward-time FTLE are found to bound vortical regions of flow, consistent with Eulerian coherent structures from the Q-criterion. The reference map also proves suitable for material surface tracking despite not explicitly tracking particles. This capability can provide valuable insights into the Lagrangian landscape of turbulent momentum transport, complementing Eulerian velocity-field analysis. The evolution of initially wall-normal material surfaces in the viscous sublayer, buffer layer, and log layer sheds light on the Reynolds stress-generating events from a Lagrangian perspective. Eliminating the need for tracking numerous particles, the reference map approach offers a convenient and promising avenue for future investigations into Lagrangian kinematics and dynamics of fluid flows.
@@ 1 Introduction
Fluid flows are often characterized by dynamic behaviors of intricate spatial patterns. Notable examples include bacteria-induced active-matter chaos (Ran et al., 2021; Urzay et al., 2017), a forest of vigorous turbulent eddies in wing boundary layers (Goc et al., 2021), and the great red spot of Jupiter (Marcus, 1993). While visually captivating, the complex interaction among these spatial patterns can sometimes obscure straightforward comprehension of the dynamically important mechanisms governing them. The concept of coherent structures has proven integral to identifying these organized features and furthering the analysis and modeling of their dynamical impact on the flow. Coherent structures, broadly defined, represent persistent patterns or organized motions that stand out against the background flow. In a more specific sense, Hussain (1986) defined coherent structures in turbulent flows as “ a connected turbulent fluid mass with instantaneously phase-correlated vorticity over its spatial extent.” Meanwhile, Haller (2015) offered a definition strongly rooted in Lagrangian/material kinematics and dynamical systems theory, describing them as “special surfaces of fluid trajectories that organize the rest of the flow into ordered patterns.” Despite differences in the perspectives and approaches taken to define them, it is perhaps unanimously agreed that coherent structures serve as the engines driving the complex interplay of momentum, heat, and scalar transport at various regimes.
@@
Recognizing the crucial role coherent structures play in fluid dynamics, the fluids community has dedicated considerable attention to developing methods for elucidating and understanding these organized features. A widely used class of Eulerian methods, particularly within the turbulence community, involves vortex identification schemes based on the invariants of the velocity gradient tensor or its symmetric/antisymmetric decomposition. The well-known $Q$-criterion (Hunt et al., 1988) extracts vortical structures from the second invariant of the velocity gradient tensor $\nabla u$, defined as $Q=-\frac{1}{2}\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{j}}{\partial x_{i}}=\frac{1}{2}(||\Omega||^{2}-||S||^{2})$, where $\Omega$ and $S$ are the rate-of-rotation and rate-of-strain tensors, respectively. $Q$ is regarded as the strength of the rotational component of the velocity field relative to the shear, and isosurfaces of $Q$ at a selected positive threshold value are used as candidate coherent structures. In response to potential inaccuracies of the $Q$-criterion when vortices are subjected to strong external strain, Jeong & Hussain (1995) proposed the $\lambda_{2}$-criterion. The $\lambda_{2}$-criterion imposes the condition that $\lambda_{2}<0$, where $\lambda_{2}$ is the median of three eigenvalues of $S^{2}+\Omega^{2}$. This requirement is based on the principles of Galilean invariance and the presence of net vorticity, which is correlated with a local pressure minimum at the vortex core. Zhou et al. (1999) proposed a detection scheme based on complex eigenvalues of $\nabla u$. From the conjugate pair of the complex eigenvalues $\lambda_{cr}\pm i\lambda_{ci}$ of $\nabla u$, the imaginary part $\lambda_{ci}$ was shown to represent the local swirling strength of the vortex. In an extension of this work, Chakraborty et al. (2005) suggested an additional constraint $\lambda_{cr}/\lambda_{ci}<\delta$ to ensure the compactness of the spiral material orbits. The aforementioned detection methods are computed from the instantaneous Eulerian velocity/tensor fields, making their application to spatial velocity data obtained from numerical simulations or experiments straightforward. They have proven effective in excluding regions of strong shear without net swirling motion. These methods necessitate users to prescribe arbitrary thresholds for isosurface detection. As demonstrated by Pierce et al. (2013), all the aforementioned schemes produce nearly identical landscapes of turbulent eddies with an appropriate selection of threshold values for each scheme.
@@
In contrast to Eulerian methods described in the previous paragraph, Lagrangian coherent structures (LCS) offer an alternative approach to coherent structure detection with an emphasis on coherence at the material (Lagrangian) level and objectivity. Haller (2005) demonstrated that the previously mentioned Eulerian diagnostics, while Galilean invariant, fail the test of objectivity, where the coherent structures identified should be invariant under Euclidean coordinate changes, including time-dependent frame rotation and translation. It’s essential to note that this does not diminish the utility of Eulerian diagnostics altogether. These approaches effectively reveal coherent patterns in the instantaneous velocity field, aiding in understanding vortex interaction/formation and momentum transport in wall-bounded flows (Jeong & Hussain, 1995; Zhou et al., 1999; Duguet et al., 2012; Motoori & Goto, 2019; Yao & Hussain, 2020; Chong et al., 1998). However, limitations may arise in the presence of frame rotation, and structural models built from the detected patterns may have constraints concerning objectivity. Haller recognized that objectivity can be ensured by basing coherent structure detection on fluid particle trajectories, which are fundamentally invariant Lagrangian features of a given flow. The representation of trajectories may vary in different reference frames, but the trajectories themselves remain invariant. The spatial patterns created by passively advected tracers then naturally embody the candidates for LCS. Haller (2015) explains LCS as “the most repelling, attracting, and shearing material surfaces that form the skeleton of Lagrangian particle dynamics.” As such, LCS is expected to have direct connections to the motion of Lagrangian fluid elements.
@@ 2 Theoretical background
First, the flow map (also referred to as deformation mapping or motion map) x = $\bm{\chi}^{t}_{t_{0}}(\textbf{X},t)$ maps a fluid particle (tracer) located at X in the reference (or initial) configuration at the initial time $t_{0}$, $\mathcal{B}(t_{0})$, to a point x in the current configuration of the continuum body at time $t$, $\mathcal{B}(t)$, (see figure 1). The deformation gradient (or Jacobian of the flow map) $\mathbf{F}^{t}_{t_{0}}(\textbf{X},t)$ is defined as
@@
which maps a material point x in the current configuration at time $t$ to its initial position X in the reference configuration. Hereinafter, this inverse flow map $\bm{\xi}$ is referred to as the reference map following Gurtin et al. (2010) . The deformation gradient $\textbf{F}^{t}_{t_{0}}(\textbf{X},t)$, which is the Jacobian of the forward motion map at the initial position X,
$\left.\textbf{F}^{t}_{t_{0}}(\textbf{X},t)\right|_{\textbf{X}=\bm{\xi}(\textbf{x},t)}=\left(\frac{\partial\bm{\xi}(\textbf{x},t)}{\partial\textbf{x}}\right)^{-1},$ (3)
can be determined from the reference map. Analogously, the inverse deformation gradient (or the Jacobian of the backward motion map at the current position x) is defined as
@@
The reference map framework was developed originally in the context of isothermal compressible and incompressible fluid-structure interaction (Kamrin et al., 2012; Valkov et al., 2015; Jain et al., 2019; Rycroft et al., 2020; Wang et al., 2022). Additionally, it has been employed in computational solid mechanics with non-trivial hyperelastic constitutive models undergoing inhomogeneous deformation (Kamrin et al., 2012). In this line of work, the Lagrangian solid equation recast in the Eulerian form, which resembles the Navier-Stokes equations, was solved using finite difference methods. The deformation gradient, determining the solid stress, was updated from the reference map. The key starting point was to identify x, the current position of material points, as the fixed grid points in Eulerian simulations or measurements. $\bm{\xi}(\textbf{x},t)$ then returns the initial position X of a Lagrangian material point that is arriving at the grid point x at time $t$. Since a material point labeled by X preserves its identity during the motion, the Lagrangian time derivative of the reference map is the zero vector,
@@
In the present study, we employ the reference map to compute the Lagrangian kinematics of a fluid, where the reference map is updated through Eq.(5) in an Eulerian manner. That is, the reference map is passively advected on the same fixed Eulerian grid utilized in fluid dynamics simulations or from experimental velocity measurements. The deformation gradient is then determined entirely from the reference map, which can be further processed to compute Lagrangian coherent structures (LCSs). Additionally, the reference map provides a natural means for material surface tracking. Details of these two tasks are provided in the following subsections. Computational details associated with the discretization of Eq.(5) and initial/boundary conditions for the reference map are provided in Section 3.
Note that an Eulerian method to compute FTLE fields utilizing Eq. (5) has been previously proposed by Leung (2011) in the context of the level set method, with application to simple flows (Leung, 2013; You & Leung, 2014, 2018a, 2018b, 2020, 2021). Here, we connect this approach to the reference map technique, and apply it to complex three-dimensional turbulent flows for the first time. Furthermore, we discretize the conservative form of Eq. (5) utilizing low dissipation numerical methods important for scale-resolving simulation of turbulent fluid flow (i.e., a second order central scheme), as compared to dissipative schemes (e.g., WENO) used by previous studies. We also note that the use of the reference map for Lagrangian material surface tracking to investigate turbulent flow structures and momentum transport in wall-bounded flows is a novelty that differs from previous applications of this technique.
@@ 2.1 Computation of Lagrangian coherent structure (LCS)
    $\Lambda^{t_{0}}_{t}(\textbf{x})=\frac{1}{|t-t_{0}|}\text{ln}\sqrt{\lambda_{\text{max}}(\textbf{C}^{t_{0}}_{t}(\textbf{x}))}.$ (11)
@@
Note that the backward-time FTLE is defined in the current configuration (x), which can be directly computed at the current position (grid points) using the reference map technique. Conversely, the forward-time FTLE is, in principle, defined in the reference configuration X, which can be directly computed using a Lagrangian particle approach. However, the reference map technique can be used to compute the image, under the forward flow map, of the FTLE field in the current configuration. This can be shown by inverting Proposition 1 from Haller & Sapsis (2011) to find
@@
which shows that the largest forward-time FTLE can be obtained from the smallest backward-time FTLE defined at the current position x. Proof of this equation is provided in Appendix A.
@@ 2.2 Lagrangian tracking of material surface
In the original work by Kamrin et al. (2012) which introduced the reference map technique for fluid/structure interaction, the reference map was used primarily for computing the solid stress tensor, and the material interface was tracked separately by integrating the level-set equation. Later, Valkov et al. (2015) observed that a level set field for the interface could be constructed directly from the reference map. Jain et al. (2019) and Rycroft et al. (2020) used this to eliminate the need for separately integrating the level set equation, observing that the reference map provides a direct mapping to the initial positions of the advected particles. This feature allows for convenient updates of the level-set function or material interface directly from the reference map. Adopting this concept, particularly from Eq. (25) in Jain et al. (2019), the reference map can be demonstrated to offer basic capabilities for Lagrangian tracking of material surfaces in purely fluid flows. For instance, in two-dimensional flows, consider a scalar field constructed from the reference map
@@
3 Numerical details
We turn our attention to the implementation of Eq. (5), the passive scalar advection equation for the reference map $\bm{\xi}$ (or strictly, its components $\xi_{1}$, $\xi_{2}$ and $\xi_{3}$, which are scalar quantities). We choose to solve the conservative form of Eq. (5) following Kamrin et al. (2012) ,
$\bm{\xi}(\textbf{x},t=t_{0})=\textbf{x}.$ (14)
In this study, we consider solid walls, symmetry planes, and periodic boundary conditions. On the boundaries where $u_{n}=0$ is expected (here $u_{n}$ is the boundary-normal velocity) such as the solid wall or symmetry planes, the advective flux of $\bm{\xi}$ can be set to zero ($\xi_{i}u_{n}=0$, $i=1,2,3$). The periodic boundary condition, which is straightforward to implement for the flow variables, is problematic for the reference map, especially for $\xi_{1}$, where $x_{1}$ corresponds to the main flow direction. When Eq. (14) is used for the initial condition, the reference map variables are linear functions of x, and periodicity cannot be enforced inherently. Advection of such discontinuous data with central schemes can be problematic, because the dispersion error may lead to the formation of wiggles and their spread along or opposite to the flow direction, as shown in figure 2 for $Re_{\tau}=180$ turbulent channel. This is often accompanied with violation of the boundedness for passively advected scalars, where the reference map at all times must be bounded by its range from the initial condition. This can be remedied by rewriting Eq. (13) in terms of the displacement map
Note that, unlike the reference map, periodicity of the displacement map is guaranteed, provided that its initial condition and the velocity field are periodic. Fortunately, from Eq. (14), the initial condition for the displacement map is a zero field which is periodic. Advection of the displacement field therefore does not suffer from the dispersion error. The strategy for periodic boundary conditions then will be to update the displacement field via Eq. (16), and to recover the reference map via Eq. (15). Note that $\bm{g}$, especially its streamwise component, will in general grow/decay in time due to the right-hand side of Eq. (16), and the reference map obtained from Eq. (15) may indicate initial particle locations outside of the computational domain. In the spirit of periodicity, this can be corrected by keeping adding/subtracting the spatial period to/from the reference map until it falls within the domain. While the reference map so obtained is free from oscillations, it does exhibit a discontinuity front that is advected in the flow, potentially impacting the computation of derived quantities, such as the FTLE field as shown in the bottom plot in figure 2. Note that it is still easy to identify extensive regions of the flow that do not contain the discontinuity.
Although not explored in the present study, on the outflow boundary zone where the flow is expected to exit normal to the boundary, one may choose to treat the boundary data as unknown and implement a one-dimensional convective outflow boundary condition in the boundary normal direction (Schlüter et al., 2005), which highly resembles the equation for $\bm{\xi}$.
Figure 3: Comparison of the FTLE computation cost on MATLAB for the 2D Taylor-Green vortex (Sec. 4.1). Blue solid line, particle-based approach (LCS Tool); red circles, asymptote for linear scaling; black solid line, reference map approach where the time step was fixed to $dt$=2e-2, corresponding to the maximum CFL condition on the finest grid; black dashed line, reference map approach with the fixed CFL number (1.6). The effect of the time step on the computed FTLE field was seen to be negligible in this test case. LCS Tool is a vectorized code, whereas the reference map code is not.